OpenStudy (anonymous):

Asked AGAIN . The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1070! If three times the number of one-step equations minus twice the number of two-step equations is equal to 110, how many one-step equations auditioned today?

OpenStudy (zzr0ck3r):

is that 1070! as in 1070 factorial?

OpenStudy (anonymous):

no its just 1070!

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20110903113658AANB9Kr go here and u will see your problem

OpenStudy (anonymous):

it should help u man!

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

your welcome!

OpenStudy (zzr0ck3r):

stop with the ! it means something in math. thats like ending everything with a +.:)

OpenStudy (anonymous):

i never knew that

OpenStudy (anonymous):

yes it means factorial for instance 8! = 8*7*6*5*4*3*2*1 and 5! = 5*4*3*2*1

OpenStudy (anonymous):

oh wow i feel dumb o.o

OpenStudy (zzr0ck3r):

we all did.